OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Two divided by x squared. + 3 \[\frac{ 2 }{x^2}+3\]

OpenStudy (anonymous):

there are lots of ways to do this

OpenStudy (anonymous):

you could say \(g(x)=x^2\) and \(f(x)=\frac{1}{x}+3\) for example

OpenStudy (anonymous):

oops no, \(f(x)=\frac{2}{x}+3\)

OpenStudy (anonymous):

or you could say \(g(x)=\frac{2}{x^2}\) and \(f(x)=x+3\)

OpenStudy (anonymous):

or \(g(x)=\frac{1}{x^2}\) and \(f(x)=2x+3\)

OpenStudy (anonymous):

i bet you can make up another one if you try

OpenStudy (anonymous):

omg thanks!